OpenStudy (toxicsugar22):

12x^2-3y^2+24y-84=0 give the exact coordinates of any four points on the hyperbola

OpenStudy (anonymous):

again?

OpenStudy (toxicsugar22):

yes ou didnt give me a good answer

OpenStudy (toxicsugar22):

it wasnt right

OpenStudy (toxicsugar22):

hi can u help me

OpenStudy (amistre64):

lets see what we can accomplish

OpenStudy (watchmath):

12x^2-3(y^2-8y)=84 12x^2-3(y-4)^2=100

OpenStudy (watchmath):

no it was wrong... forget it :)

OpenStudy (amistre64):

(12x^2) + (-3y^2+24y) -84=0 (12x^2) +-3 (y^2-8y + ___) =84 +-3(____) (12x^2) +-3 (y^2-8y + 16) =84 +-3(16) 12x^2 -3 (y-4)^2 =84 -48 = 36 12x^2 3(y-4)^2 ----- - -------- = 1 36 36 (1/3)x^2 - (1/12)(y-4^2) = 1

OpenStudy (toxicsugar22):

ok

OpenStudy (amistre64):

when y = 4; x = +- sqrt(3) so.. (+sqrt(3),4) (-sqrt(3),4) are 2 points we know of; they are the vertices of the hyperola right?

OpenStudy (toxicsugar22):

ok

OpenStudy (amistre64):

when x=3; we get 3-y = 1 y will have to even out to 2 then... (y-4)^2 ------ = 2 12 (y-4)^2 = 24 y^2 -8y +16-24 = 0 y^2 -8y -8 = 0 y = -4 +- 2sqrt(2) can you input a sum of numbers for a point?

OpenStudy (toxicsugar22):

i dk

OpenStudy (amistre64):

i might be over thinking that; lets try at y = 0 (1/3)x^2 - 16/12 = 1 x^2 = (12+16/4) x^2 = 7 x = +- sqrt(7) then is simpler (sqrt(7),0) (-sqrt(7),0) are another 2 points

OpenStudy (anonymous):

hi. didn't we get exactly this answer before? was it wrong? because it looks good to me.

OpenStudy (amistre64):

sqrt(3) need to dbl chk it

OpenStudy (amistre64):

36 -48+96 - 84 ?=0

OpenStudy (amistre64):

132 -132 = 0 is good

OpenStudy (amistre64):

sqrt(7) is good cause 12.7 =84

OpenStudy (toxicsugar22):

good thanks

OpenStudy (toxicsugar22):

can u help me with one more

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0

OpenStudy (toxicsugar22):

give the exact coordinated of an four points

OpenStudy (toxicsugar22):

can u helpme